3. (25 points) Solve the following differential equations by using variation of parameters. y + y = sec x -
Answer #1

(62+) y = secx. This is of the form f(D)-Y = 2(X). Wher FID)=D²+1 Q(X) = secz. Auriliary equation as m+1=0. ie firm)=0 mati